# From Images to Insights: Explainable Biodiversity Monitoring with Plain Language Habitat Explanations

### v1

@@ Abstract
Explaining why the species lives at a particular location is important for understanding ecological systems and conserving biodiversity. However, existing ecological workflows are fragmented and often inaccessible to non-specialists. We propose an end-to-end visual-to-causal framework that transforms a species image into interpretable causal insights about its habitat preference. The system integrates species recognition, global occurrence retrieval, pseudo-absence sampling, and climate data extraction. We then discover causal structures among environmental features and estimate their influence on species occurrence using modern causal inference methods. Finally, we generate statistically grounded, human-readable causal explanations from structured templates and large language models. We demonstrate the framework on a bee and a flower species and report early results as part of an ongoing project, showing the potential of the multimodal AI assistant backed up by a recommended ecological modeling practice for describing species habitat in human-understandable language.
@@ 1 Introduction
Given a photo, our system automated achieves: (1) identify the species, (2) locate global occurrences via biodiversity APIs, (3) simulate absence background conditions, (4) extract environmental data, (5) learn the causal structure among climatic variables, (6) infer their causal effect on species occurrence, and (7) explain the findings in nature language.
@@ 2 Related Works
Visual Species Recognition. Existing works in species recognition have achieved high performance. The most notable is BioCLIP[18], a foundation model trained on 10 million images. Recently, BioCLIP 2[11] incorporates hierarchical taxonomic embeddings and demonstrates ecological semantic alignment with ecological traits. The iNaturalist Challenge dataset is designed to benchmark the classification and detection of fine-grained species in the wild. Balk et al.[4] present an automated workflow to extract biological knowledge in the emerging field of imageomics from images.
Species Distribution Modeling (SDM). Species distribution models (SDMs) are defined as empirical models that aim to estimate the probability of species distributions across spatial and temporal dimensions[20]. Pseudo-absence generation techniques are proposed to simulate true absences when unavailable. Barbet-Massin et al.[5] explore several strategies for selecting pseudo-absences. Deep learning-based species distribution models (deep SDMs)[8] use geographic predictors and citizen-science data to address sampling gaps.
Causal Graph Learning. Causal discovery aims to identify directed relationships among variables from data. NOTEARS[22] introduces a continuous optimization framework for learning Directed Acyclic Graph (DAGs). DAG-GNN[21] extends the framework to nonlinear and variational settings. Santos et al.[16] conduct a systematic literature review with data-driven causal inference analysis. Luo et al.[13] develop a causal discovery model that effectively identifies the causal network within the ecosystem.
@@ 3.2 Occurrence Data Retrieval
Given the predicted species name, we query the Global Biodiversity Information Facility (GBIF)111GBIF — Global Biodiversity Information Facility, the largest occurrence database with over 3.1 billion records (accessed June 2025). We retrieve all occurrence records with the following filters:
@@ 3.3 Pseudo-Absence Sampling
True absence data is rarely available in ecological datasets. To construct a binary classification framework (Presence 0,1), we improve the pseudo-absence sampling strategy inspired by [5]: (1) Expand the bounding box by ±1° latitude and longitude (111 km): Create a spatial buffered region around occurrences to define the sampling area. (2) Equally sample 2 points within the buffered region: Randomly generate background points without spatial bias. (3) Filter out points within 5 km of a presence record: To avoid mislabeling the likely presence of species as absence. (4) Mask out marine areas using Natural Earth land polygons222Natural Earth — Land 1:10m: Ensure that background points fall only on terrestrial land. These pseudo-absence points are treated as negative samples (Presence = 0).
@@ 3.7 Explanation Generation
Finally, we generate causal explanations using two complementary approaches: (1) Rule-based generation, following the template ( “BIO” is bioclimatic variable name, “SP.” is species name), such as, If ATE 0 (ATE 0.1)“High BIO strongly promotes SP. presence. This likely reflects a core habitat requirement.”; If ATE 0 (-0.05 ATE 0) “BIO has a weak negative effect.”. (2) LLM-enhanced generation, using LLAMA3.3-70B[14] with constrained prompting. To ensure scientific rigor, prompts are designed to ensure factual grounding (such as “Ensure the explanation is realistic, grounded in ecological reasoning, and free from vague generalizations”, “Write 1 sentence explaining the most likely ecological mechanism behind the causal influence”). The output is designed for environmental scientists, conservation policy experts, or education platforms requiring accessible scientific communication.
@@ 4.1 Pilot Study Setup
To assess the effectiveness and interpretability of our framework, we conduct experiments on two representative species from different ecological niches as pilot studies: Osmia parietina (Western mason bee), a wild European bee that nests in sunny mountain habitats, and Ajuga reptans (bugle), a ground-covering blue spring flower that thrives in moist conditions. Data retrieved via GBIF for each species is 1,000 presence points, with 2,000 background points. The final data structure is: [Latitude, Longitude, BIO1, …, BIO19, Presence (0/1)].
@@ 4.2 Causal Analysis and Example Explanations
An example of Ajuga reptans, the learned environmental DAGs revealed strong dependencies, such as BIO 11 (Mean Temperature of Coldest Quarter) → BIO6 (Min Temperature of Coldest Month), BIO 10 (Mean Temperature of Warmest Quarter) → BIO7 (Temperature Annual Range). Then DoWhy estimated statistically ATEs = 0.13 for BIO11 and ATEs = -0.03 for BIO10. This suggests that its presence is positively influenced by milder winters. However, high summer temperatures, extreme temperature fluctuations limit its distribution and abundance.
@@ 5 Conclusion
We introduced a novel visual-to-causal, end-to-end framework that transforms a species image into interpretable causal insights about its habitat in plain language. Our approach integrates species recognition, geospatial data retrieval, climate variable extraction, causal graph discovery, and causal effect estimation into a single automated pipeline. This provides a novel perspective to answer why a species is found where it is, grounding predictions in interpretable ecological mechanisms rather than black-box patterns.
Although the model currently captures only a limited subset of relevant variables, it shows significant potential in conservation planning, climate risk assessment, and ecological education, and provides a foundation for future AI systems that reason about the natural world in transparent and human-aligned ways.
@@ 5.0.1 Acknowledgements
The project on which this report is based was funded by the German Federal Ministry of Education and Research within the Research Initiative for the Conservation of Biodiversity (FEdA) under the funding code 16LW0682K. The responsibility for the content of this publication lies with the author.

### v2

@@ Abstract
Explaining why the species lives at a particular location is important for understanding ecological systems and conserving biodiversity. However, existing ecological workflows are fragmented and often inaccessible to non-specialists. We propose an end-to-end visual-to-causal framework that transforms a species image into interpretable causal insights about its habitat preference. The system integrates species recognition, global occurrence retrieval, pseudo-absence sampling, and climate data extraction. We then discover causal structures among environmental features and estimate their influence on species occurrence using modern causal inference methods. Finally, we generate statistically grounded, human-readable causal explanations from structured templates and large language models. We demonstrate the framework on a bee and a flower species and report early results as part of an ongoing project, showing the potential of the multimodal AI assistant backed up by a recommended ecological modeling practice for describing species habitat in human-understandable language. Our code is available at https://github.com/Yutong-Zhou-cv/BioX.
@@ 1 Introduction
Given a photo, our system automated achieves: (1) identify the species, (2) locate global occurrences via biodiversity APIs, (3) simulate absence background conditions, (4) extract environmental data, (5) learn the causal structure among climatic variables, (6) infer their causal effect on species occurrence, and (7) explain the findings in nature language. Our contribution extends beyond combining existing models, providing a novel multimodal perspective for explainable species distribution modeling and ecological reasoning.
@@ 2 Related Works
Visual Species Recognition. Existing works in species recognition have achieved high performance. The most notable is BioCLIP[20]. Recently, BioCLIP 2[13] incorporates hierarchical taxonomic embeddings and demonstrates ecological semantic alignment with ecological traits. The iNaturalist Challenge dataset is designed to benchmark the classification and detection of fine-grained species in the wild. Balk et al.[4] present an automated workflow to extract biological knowledge in the emerging field of imageomics from images.
Species Distribution Modeling (SDM). Species distribution models (SDMs) are defined as empirical models that aim to estimate the probability of species distributions across spatial and temporal dimensions[22]. Our framework shifts the focus from correlation-based prediction to causally grounded and interpretable insights. Pseudo-absence generation techniques are proposed to simulate true absences when unavailable. Barbet-Massin et al.[5] explore several strategies for selecting pseudo-absences.
Causal Graph Learning. Causal discovery aims to identify directed relationships among variables from data. NOTEARS[24] introduces a continuous optimization framework for learning Directed Acyclic Graph (DAGs). DAG-GNN[23] extends the framework to nonlinear and variational settings. Santos et al.[18] conduct a systematic literature review with data-driven causal inference analysis.
@@ 3.2 Occurrence Data Retrieval
Given the predicted species name, we query the Global Biodiversity Information Facility (GBIF)111GBIF — Global Biodiversity Information Facility: https://www.gbif.org/, the largest occurrence database with over 3.1 billion records (accessed June 2025). We retrieve all occurrence records with the following filters:
@@ 3.3 Pseudo-Absence Sampling
True absence data is rarely available in ecological datasets. To construct a binary classification framework (Presence 0,1), we improve the pseudo-absence sampling strategy inspired by [5]: (1) Expand the bounding box by ±1° latitude and longitude (111 km): Create a spatial buffered region around occurrences to define the sampling area. (2) Equally sample 2 points within the buffered region: Randomly generate background points without spatial bias. (3) Filter out points within 5 km of a presence record: To avoid mislabeling the likely presence of species as absence. (4) Mask out marine areas using Natural Earth land polygons222Natural Earth — Land 1:10m: https://www.naturalearthdata.com/downloads/10m-physical-vectors/10m-land/: Ensure that background points fall only on terrestrial land. These pseudo-absence points are treated as negative samples (Presence = 0).
@@ 3.7 Explanation Generation
Finally, we generate causal explanations using two complementary approaches: (1) Rule-based generation, following the template ( “BIO” is bioclimatic variable name, “SP.” is species name), such as, If ATE 0.1, generate “High BIO strongly promotes SP. presence. This likely reflects a core habitat requirement.”; If -0.05 ATE 0, generate “BIO has a weak negative effect.”. (2) LLM-enhanced generation, using LLAMA3.3-70B[16] with constrained prompting. To ensure scientific rigor, prompts are designed to ensure factual grounding (such as “Ensure the explanation is realistic, grounded in ecological reasoning, and free from vague generalizations”, “Write 1 sentence explaining the most likely ecological mechanism behind the causal influence”). The output is designed for environmental scientists, conservation policy experts, or education platforms requiring accessible scientific communication.
@@ 4.1 Pilot Study Setup
To assess the effectiveness and interpretability of our framework, we conduct experiments on two representative species from different ecological niches as pilot studies: Osmia parietina (Western mason bee), a wild European bee that nests in sunny mountain habitats, and Ajuga reptans (bugle), a ground-covering blue spring flower that thrives in moist conditions. Data retrieved via GBIF for each species[11, 12] is 1,000 presence points, with 2,000 background points. The final data structure is: [Latitude, Longitude, BIO1, …, BIO19, Presence (0/1)].
@@ 4.2 Causal Analysis and Example Explanations
An example of Ajuga reptans, the learned environmental DAGs revealed strong dependencies, such as BIO 11 (Mean Temperature of Coldest Quarter) → BIO6 (Min Temperature of Coldest Month), BIO 10 (Mean Temperature of Warmest Quarter) → BIO7 (Temperature Annual Range). Then DoWhy estimated statistically ATEs = 0.13 for BIO11 and ATEs = -0.03 for BIO10. This suggests that its presence is positively influenced by milder winters. However, extreme temperature fluctuations limit its distribution and abundance.
@@ 5 Conclusion
We introduced a novel visual-to-causal end-to-end framework that transforms species images into causal ecological explanations. By integrating species recognition, geospatial retrieval, climate variable extraction, causal graph learning, and inference, the pipeline provides an interpretable alternative to correlation-based distribution modeling and answers why species occur where they do.
The study remains preliminary, as only two case studies were presented, and the ecological validity of the explanations has not yet been systematically assessed. Despite these limitations, the approach shows strong potential for conservation planning, climate risk assessment, and ecological education, and provides a foundation for future AI systems that reason about the natural world in transparent and human-aligned ways.
@@ 5.0.1 Acknowledgements
The project on which this report is based was funded by the German Federal Ministry of Education and Research within the Research Initiative for the Conservation of Biodiversity (FEdA) under the funding code 16LW0682K. The responsibility for the content of this publication lies with the author. We thank the data publishers who made their datasets available through GBIF.org and whose contributions are listed on the respective download DOI pages.